The one-year interest rate is 4% in 2019. The two year interest rate is 5% 2019 and the

three-year interest rate is 6% in 2019. According to the expectation hypothesis, what is

your expectation of one-year interest rate in 2020 and 2021 respectively? Explain your

calculation.
Answer #1

Answer

According to Expectation hypothesis :

Rnt = (R1t + R1t + 1 + R1t + 1 + ---------------R1t + n - 1)/n

where, Rnt = n periods interest rate and R1t + n - 1 = one year expected interest rate of nth period.

Now, Here The one-year interest rate is 4% in 2019. The two year interest rate is 5%

i.e. R2t = 2 year interest rate = 5%, R1t = one year interest rate in 2019 = 4% and R1t + 1 = one-year interest rate in 2020 is what we have to calculate.

Using above formula we have :

R2t = (R1t + R1t + 1)/2

=> 5% = (4% + R1t + 1)/2

=> R1t + 1 = 6%

Hence, expectation of one-year interest rate in 2020 = 6%

Now, Here The one-year interest rate is 4% in 2019. The three year interest rate is 6% 2019

i.e. R3t = 3 year interest rate = 5%, R1t = one year interest rate in 2019 = 4%, R1t + 1 = one year interest rate in 2020 = 6% , R1t + 2 = one-year interest rate in 2021 is what we have to calculate.

Using above formula we have :

R3t = (R1t + R1t + 1 + R1t + 2)/3

=> 6% = (4% + 6% + R1t + 2)/3

=> R1t + 1 = 18% - 4% - 6% = 8%

Hence, expectation of one-year interest rate in 2021 = 8%